Sargent Corporation applies overhead cost to jobs on the basis of 90% of direct labor cost. If Job 210 shows $22,860 of manufacturing overhead cost applied, how much was the direct labor cost on the job? Multiple Choice $25,400 $43,434 $20,574 $22,860
Answer #1
A For every $1 of direct labor cost, Applied overhead = 0.9$
B For 0.9 $ manufacturing overhead applied, Direct labor cost $1
C=B/A For $1 manufacturing overhead overhead applied , direct labor cost $1.11 (1/0.9)
D Manufacturing overhead applied $22,860
E=C*D For $22860 manufacturing overhead applied, Direct labor cost= $        25,400
F Direct Labor Cost $        25,400
G=0.9*F Manufacturing overhead applied $        22,860
Answer:
$                                                                                                                             25,400